The Munder Difflin Paper Corporation provides an executive stock option plan. Under the plan, the company granted options to the CEO on January 1, 2013, that permit her to acquire 12 million of the company's $1 par value common shares within the next five years, but not before December 31, 2014 (the vesting date). The exercise price is the market price of the shares on the date of the grant, $22 per share. The fair value of the options, estimated by an appropriate model, is $4 per option.

On April 30, 2016, the market price of the common stock is $50 so the CEO exercised 9 million options to purchase 9 million shares.

a) Write the compensation expense that will be recorded in 2013 and 2014, show work

b) Prepare the appropriate journal entry on April 30, 2016 to record this transaction, show work
Answer #1

Here ,

Stock grant = 12000000 nos

Vesting period = 1st Jan 2013 to 31st Dec 2014 = 2 Years

fair value of options =$4 per share

Upto 2013 Service period =1 year

Upto 2014 Service period = 2 years

Hence Stock options compensation price = 12000000 x $4 =$48,000,000

Compensation expenses on 2013 = $48,000,000 x service period /vesting period

= $48,000,000 x 1 /3=$16,000,000

Compensation expenses on 2014 = $48,000,000 x 2 /3= $32,000,000.

Journal entry :-

Debit Credit
Cash A/c $36,000,000
Employee Stock Options Outstanding A/c $198,000,000
Paid up Equity $9,000,000
Share premium accounts $225,000,000

Note : Cash = 9000000 x Option price = 9000000 x $4 =$36,000,000

Employee Stock Options Outstanding A/c = 9000000 x Grant price = 9000000 x $22=$198,000,000

Paid up Equity = 9000000 x fair value = 9000000 x $1 =$9,000,000

Share premium accounts = ($36,000,000 + $198,000,000-$9,000,000)=$225,000,000